Circular No. 8261

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2003le–2003ln M. Hicken, Harvard-Smithsonian Center for Astrophysics, on behalf

of the ESSENCE project (cf. IAUC 8251, 8237, 8034, 8021), reports the discovery of ten supernovae on multiple R- and I-band images taken at the Cerro Tololo Blanco 4-m telescope during 2003 Dec. 13–29 by project members. Spectroscopy was obtained on Dec. 13–29 with the Gemini North and South telescopes (+ GMOS), the MMT (+ Blue Channel), and the Keck I telescope (+ LRIS), showing each to be of type Ia, with the exception of SN 2003lg, which is of type II. Each new supernova is within 1′′ of its host-galaxy center, except for 2003lg, which is 4′′.5 away.

SN  2003 UT  α_{2000}  δ_{2000}  R  z
2003le Dec. 13 1^{h}08^{m}08.^{s}73 + 0^{o}27′09′′.7 22.7 0.6
2003lf Dec. 13 1 08 49.81 − 0 44 13.5 22.1 0.4
2003lg Dec. 13 1 13 32.68 − 0 36 57.3 22.5 0.2
2003lh Dec. 13 2 10 19.51 − 4 59 32.3 22.7 0.5
2003li Dec. 13 2 27 47.29 − 7 33 46.2 22.8 0.5
2003lj Dec. 15 1 12 10.03 + 0 19 51.3 22.1 0.4
2003lk Dec. 15 2 11 12.82 − 4 13 52.1 22.8 0.4
2003ll Dec. 15 2 35 41.19 − 8 06 29.6 23.2 0.6
2003lm Dec. 15 23 24 25.51 − 8 45 51.1 22.5 0.4
2003ln Dec. 15 23 30 27.15 − 8 35 47.0 22.9 0.6

SUPERNOVAE 2003lo AND 2003lp Further to IAUC 8260, T. Puckett reports the discovery of two appar-

ent supernovae on unfiltered CCD frames taken with his 0.50-m automated supernova patrol telescope — SN 2003lo in collaboration with B. Kerns (Stanton, ND) and SN 2003lp with D. Toth (Lima, OH).

SN  2003 UT  α_{2000}  δ_{2000}  Mag.  Offset
2003lo Dec. 31.14 3^{h}37^{m}05.^{s}12 − 5^{o}02′17′′.3 17.2 11′′.2 W, 16′′.3 N
2003lp Dec. 31.46 11 44 28.57 +69 43 54.6 16.9 6′′.9 W, 7′′ N

SN 2003lo in NGC 1376 was also present on unfiltered CCD frames taken on 2004 Jan. 1.15 UT, and SN 2003lp in UGC 6711 was also present on frames taken on 2004 Jan. 1.23. Neither apparent supernova was present on images taken by Puckett during 1998–2002.

2004 January 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green